1. An object of mass 20 kg is pushed on a floor with a force of 40sin 2t N. Given that the 

frictional force is 20 times the velocity and the object starts from rest, determine the 

velocity of the object as a function of time. 


2. Consider an LCR circuit with L = 0.1 H, C = 0.01 F and R = 3.0. Determine the 

electric current in the circuit, given that at t = 0, the charge in the circuit is zero and the 

current is 2 A.




1
Expert's answer
2021-05-19T12:52:53-0400

(1) Force applied, "F=40sin(2t)"

Let force of friction "=F_f=m\dfrac{d(20v)}{dt}=20m\dfrac{dv}{dt}"

Net force, "F_{net}=F-F_f=40sin(2t)-20m\dfrac{dv}{dt}"

"ma=40sin(2t)-20m\dfrac{dv}{dt}"

"a=\dfrac{40sin(2t)}{m}-20\dfrac{dv}{dt}"

"\dfrac{dv}{dt}=\dfrac{40sin(2t)}{m}-20\dfrac{dv}{dt}"

"21dv=\dfrac{(40sin(2t))}{m}dt"

"\intop^v_odv=\dfrac{1}{21}\intop^t_o\dfrac{(40sin(2t))}{m}dt"

"v=\dfrac{40}{21\times20\times2}[-cos(2t)+1]"

"v=0.047(1-cos2t)"


(2) Inductance, "L=0.1\space H"

Capacitance, "C=0.01\space F"

Resistance, "R=3\space \Omega"

By Kirchoff's law we have,

"L\dfrac{d^2q}{dt^2}+R\dfrac{dq}{dt}+\dfrac{q}{C}=0"


taking "R/L=2\lambda=30" and "1/LC=\mu^2=1000" we get

"\lambda=15" and "\mu=31.622"

"\dfrac{d^2q}{dt^2}+2\lambda\dfrac{dq}{dt}+\mu^2q=0"


"D^2+2\lambda D+\mu^2=0"

"D^2+30D+1000=0"


Since, "\mu>\lambda" , the roots of the auxiliary equation are imaginary, i.e,

"D=-\lambda\pm i\alpha"

where, "\alpha^2=\mu^2-\lambda^2"


The solution of the equation is "q=e^{-\lambda t}(c_1\space cos\space \alpha t+c_2\space sin\space \alpha t)"


at t=0, "q=0"

"0=e^0(c_1+0)"

"c_1=0"


differentiating q with respect to time,

"\dfrac{dq}{dt}=-\lambda e^{-\lambda t}(c_1\space cos\space \alpha t+c_2\space sin\space \alpha t)+e^{-\lambda t}\space\alpha(-c_1\space sin\space \alpha t+c_2\space cos\space \alpha t)"

"i(t)=-\lambda e^{-\lambda t}(c_1\space cos\space \alpha t+c_2\space sin\space \alpha t)+e^{-\lambda t}\space\alpha(-c_1\space sin\space \alpha t+c_2\space cos\space \alpha t)"


At t=0, i=2

"2=\alpha (c_2)"

"c_2=\dfrac{2}{\alpha}=\dfrac{2}{27.83}=0.07"


Substituting the values in the above equation,

"i(t)=c_2e^{-\lambda t}(\alpha\cos\alpha t-\lambda\sin\alpha t)"

"i(t)=0.07e^{-15t}(27.83\cos(27.83t)-15\sin(27.83t))"